# Generating Synthetic Free-text Medical Records with Low Re-identification Risk using Masked Language Modeling

### v1

@@ Abstract
In this paper, we present a system that generates synthetic free-text medical records, such as discharge summaries, admission notes and doctor correspondences, using Masked Language Modeling (MLM). Our system is designed to preserve the critical information of the records while introducing significant diversity and minimizing re-identification risk. The system incorporates a de-identification component that uses Philter to mask Protected Health Information (PHI), followed by a Medical Entity Recognition (NER) model to retain key medical information. We explore various masking ratios and mask-filling techniques to balance the trade-off between diversity and fidelity in the synthetic outputs without affecting overall readability. Our results demonstrate that the system can produce high-quality synthetic data with significant diversity while achieving a HIPAA-compliant PHI recall rate of 0.96 and a low re-identification risk of 0.035. Furthermore, downstream evaluations using a NER task reveal that the synthetic data can be effectively used to train models with performance comparable to those trained on real data. The flexibility of the system allows it to be adapted for specific use cases, making it a valuable tool for privacy-preserving data generation in medical research and healthcare applications.
@@ I Introduction
Among numerous synthetic data generation techniques, Casual Language Modeling (CLM) is the one currently recognised as the state-of-the-art. Encoder-decoder transformers and GPTs have showcased their capabilities to capture the statistical properties of real discharge summaries with moderate privacy risks. However, to our knowledge, no work has tried to use Masked Language Modeling for this task. In our previous work [12], we introduced the first system to generate synthetic data using Masked Language Modeling (MLM) and compared it to Casual Language Modeling at multiple synthetic generations tasks to conclude that MLM was able to outperform the latter at some synthetic data generation tasks while preserving CLM’s main ability to infer complex knowledge and relations.
@@
This paper aims to introduce a system specifically designed to generate synthetic free-text medical reports, including discharge summaries, admission notes and correspondences between doctors using Masked Language Modeling. This system includes a de-identification phase using a state-of-the-art de-identification tool called Philter [13], which does not rely on deep learning and is specifically designed for the recognition of protected health information (PHI). While most other works ignore de-identification or assume that input data are already manually labelled for de-identification, we aim to include de-identification directly within our system. Additionally, our system includes a Medical Entity Recognition model to select and retain critical medical information from the original medical records. The final system is capable of taking raw, neither annotated nor de-identified letters, and generates a large amount of unique synthetic letters that retain the critical medical information of the original letter and for which users can control its resemblance/diversity trade-off. Finally, it guarantees the reduction in the number of patient health identifiers to nearly zero at the cost of slightly reduced data fidelity. The system, training, and evaluation codes are made publicly available at https://github.com/SamySam0/SynDeidMLM.
@@ II Related Work
Using a similar system, [15] produced synthetic unstructured free-text medical data with low re-identification risk. However, the authors experienced high similarities between generated and original data, leading to reduced diversity in the synthetic data and limiting its potential in large-scale generations for uses such as oversampling. Additionally, they did not focus on de-identification pre-generation and assumed that synthetic data generation reduces re-identification risk by creating new patient records with similar, but different content. As a result, they stated that synthetic data produced must undergo rigorous de-identification of PHI elements before they can be distributed for public use.
@@
Finally, in one of the latest works on medical synthetic data generation, [18] analyzed whether GPT-3.5 could generate high quality discharge summaries. They observed that GPT-3.5 was likely to reproduce and list all concepts mentioned in the prompt. In addition, GPT-3.5 generated unnatural texts which missed necessary supporting information and tended to introduce spurious information. These are due to the hard controllability of autoregressive CLM models as to how much is kept or how much is removed, resulting in a lack of diversity in the synthetic data. That was indeed validated by clinician-evaluators who noted “the correctness of the generated discharge summaries while suffering in variety, supporting information, and coherence in narratives” (hallucinations). Finally, they revealed that GPT-3.5 failed to present diagnoses as interconnected events, which is something that may be improved through the selection of tokens to be masked when using MLM approaches.
@@ III-A The Masker
Medical Entity Recognition. The second phase in the Masker is to detect medical entities that we believe are critical information which build the core of the letter and should not be masked to preserve the letter’s intention. For this purpose, we leverage a pre-trained Medical Named Entity Recognition (NER) model which was fine-tuned on i2b2-2010 by Stanza [^1] https://stanfordnlp.github.io/stanza/available_biomed_models.html. This NER extracts 3 categories of medical entities, namely PROBLEM, TEST, and TREATMENT; and reported a micro-averaged F1 score of 88.13% on the same dataset. We use this tool to tag and protect every remaining entity in our original letter that belongs to any of these medical categories, such that they can remain in the structure of the masked letter and guide the synthetic generation. Note that the NER can be replaced at need to select different types of entities and critical information. Alternatively, a ratio can be set up to control how much of each medical entity’s category should be kept.
Part-of-Speech Tagger. The third and final phase in the Masker is to detect part of speech within the original letter. To this end, we use Stanza’s POS which assigns a part-of-speech to each remaining word based on both its definition and its context. Given the POS of all remaining entities, we select and mask a random subset of tagged entities in order to create diversity in the synthetic generations. To do so, we define a set of ratios, with values between 0 and 1 for each POS category, where values closer to 0 imply less masking and ones closer to 1 higher masking rates. For example, we could give the set {NOUN: 0.7, VRB: 0.5} to approximately mask 70% of nouns, 50% of verbs and none of the other POS categories. We leave the choice of masking ratios to the user based on their own use cases. Results of experiments on different presets are given in section V.
@@ III-B The Mask-Filling System
MLM Model. The MLM model is a deep language model which takes in the masked letter and provides a probability distribution over all possible words to replace the masked entity. As our MLM model, we selected a fine-tuned instance of the famous pre-trained encoder BERT, named Bio_ClinicalBERT [^2] https://huggingface.co/emilyalsentzer/Bio_ClinicalBERT. This model was initialized from BioBERT[20] and fine-tuned on all notes from MIMIC III [21], a database containing electronic health records from ICU patients at the Beth Israel Hospital in Boston, MA. This makes it an excellent baseline trained on both biomedical and clinical data. We then further trained this model on our specific task using the 790 training letters provided by our dataset to obtain the final version of our MLM model. Please note that we did not try alternative models as a baseline for our system. However, we truly encourage further studies to experiment with different baseline models.
@@
  • Simultaneous Chunk Filling: The first filling method takes a chunk of the masked letter and passes it to the MLM model. The latter sends back the probabilities for each masked entity. For each masked entity, the Mask-Filling component selects which word to replace it with based on the probability distribution. The user can decide to pick the word with the highest probability deterministically, or, in order to motivate diversity, he can choose to select each word stochastically from the distribution at the cost of a slightly reduced fidelity.
@@ IV-B Hyperparameter tuning and Training
During training, we perform a grid search to select the most optimal set of hyperparameters from the following values: $\alpha\in\{1\times 10^{-4},5\times 10^{-5},3\times 10^{-5}\}$, $\beta\in\{8,16\}$, $\phi\in\{0.75,1.0\}$ and $\psi\in\{0.30,0.50\}$; where $\alpha$ is the learning rate of the MLM model, $\beta$ is the training batch size, $\phi$ is the PHI’s masking proportion and $\psi$ is the overall masking probability. For convenience, we select the optimal number of training epochs through early stoppage with a patience of $p=2$. While we agree that more advanced hyperparameter search methods exist, such as Bayesian Optimization [25] or Optuna [26], we decided to opt for grid search due to computational limitations.
We split the dataset into 80% training and 20% validation, using a random split. We once again recognize that k-fold cross-validation is usually more accurate, but are constrained by the same computational resources. For each possible set of hyperparameters, a new instance of the system is created. Then, during its training, training samples are re-processed at each epoch with a random masking of up to $\psi$ percent, including $\phi$ percent of all PHI entities. This is made to allow the model to see varied versions of the same sample, increasing the diversity of cases it can learn from and thus reducing overfitting. In contrast, the validation set is masked consistently across all epochs to ensure fair comparison.
@@ IV-C System instances
  • System_S_0.7: Similarly to System_S_0.5, this instance masks all PHI entities and none of the medical entities captured by the NER. However, it masks 70% of NOUNS, VERBS and ADJECTIVES for increased diversity, and uses the same Simultaneous Chunk Filling algorithm.
@@ IV-D Evaluation Metrics
Data privacy evaluation is crucial when considering the sharing of synthetic medical data. As our current dataset has been labelled by multiple professionals following the official HIPAA-PHI de-identification rules, we evaluate the privacy level of our model by calculating the F1-score to how much of the PHIs were identified and replaced by our system according to the annotated data, and how much re-identification occurred on average post generation.
@@ V Results and Discussion
In this section, we present the results of all system instances at the three evaluation phases described in section IV. To read full synthetic samples, please refer to Appendix Examples of generated synthetic letters.
@@ V-A Lexical Similarity Evaluation against References
When looking specifically at the results of the System_S instances, increasing the masking ratios penalises the similarity more than with the iterative systems. In fact, on the same ratios (0.7), all ROUGE scores are about 3 points lower and the BERTScore 0.32 lower. That showcases one of the advantages of Iterative Mask Filling being that each mask token to fill is surrounded by original or predicted tokens, which adds to the MLM context and reduces overall uncertainty. On the other hand, when increasing System_Is’ masking ratios from 0.7 to 0.9, the ROUGE scores decrease between 3 and 5 points, while the BERTScore only decreases by less than half a point. This shows that while the generated letter is lexically further away from the original letter, the meaning is mostly preserved.
Finally, Table II presents the overlap between the top 5, 10, 20, 50 and 100 features identified across the real and synthetic datasets, without stopwords, to further evaluate the distance of the synthetic data to the original one. The results confirm that increasing the masking ratios and utilising the iterative filling algorithms further motivates diverse generations while conserving an acceptable amount of core information.
@@ V-C1 Downstream NER Task
Before starting, we split the testing set into training and testing subsets. Then the downstream pipeline is presented as follows. First, the original letters are passed through the system to generate synthetic letters. Then, the original letters and synthetic letters are passed through SciSpacy [^5] https://allenai.github.io/scispacy/ en_ner_bc5cdr_md, a SpaCy [^6] https://spacy.io/ NER model trained on the BC5CDR corpus to recognize DISEASE and CHEMICAL entities, which achieved an F1-score of 0.843 on their dataset. Note that there exist alternative NER models which could be used here, something that could be worth exploring in future works. After extraction, we obtain the detected medical entities from both the original and synthetic letters. These entities are then stored to act as an original and synthetic entities dataset to train a Spacy NER model from scratch. That is, we train a first version of the Spacy NER model on the medical NER data extracted from the original training subset, and a second version on the ones extracted from the synthetic training subset. Finally, we evaluate both versions of Spacy on the testing subset and compare their performance.
Moreover, in order to analyse how much improvement can be obtained by further augmenting the synthetic dataset, we repeat this experiment and generate two synthetic letters per original letter.
@@ V-C2 Results
The results of the downstream task are given in Table IV. As they demonstrate, all systems could achieve scores comparable to those obtained on the original letters. An interesting observation is that, the more we increase the masking ratios, the higher scores tend to be. That is probably due to the increased diversity generated in synthetic samples, allowing to generate more diverse samples for Spacy to train from. Additionally, it is interesting that the iterative mask-filling technique could achieve the highest score by a slight margin, which is in accordance with the linguistic results obtained above.
@@
TABLE IV: Average Precision, Recall and F1-score for two labels (DISEASE and CHEMICAL) using Synthetic data ×$\times 1$, ×$\times 2$ and Real data on the testing dataset.
@@ V-D Data Privacy Evaluation
Secondly, we calculated the re-identification risk, i.e. the probability that our MLM model re-introduces a valid protected health information masked by the Masker. To evaluate the re-identification risk, we leverage the risk of presence disclosure which assesses an attacker’s ability to determine if any real patient records in their possession were used to train the MLM model by comparing these records against the synthetic patient dataset. Assessing the risk of presence disclosure ensures the privacy of individuals whose data was used to train a decision model, as well as the interests of the healthcare entity where the individual received treatment. Thus, presence disclosure is a widely evaluated measure of re-identification risk. As a result, we observed that our MLM model re-identified masked PHI entities of more than two tokens with a probability of only 0.035. To give further insights into the re-identification risk, we calculated the rate of longest common substrings with lengths of 3 or more, 5 or more and 7 or more, between original and synthetic PHIs. The outcomes showed rates of only 0.098, 0.020 and 0.009, respectively.
@@ VI-A Limitations
Upon careful analysis of the generated samples, we observed that consistently filling temporal information throughout the synthetic letter, and aligning it with the original data, presents a significant challenge. Furthermore, other interconnected events such as two names within a discussion may be complicated to assign consistently when one is not present in the context’s window at the moment of generation. This is an issue that our system may encounter and is rather important as medical data inherently consists of time series, where interconnected and temporal information are critical for numerous applications, such as modelling the progression of diseases. As a result, we recommend using our system as a means to generate synthetic data used to enhance machine learning performance, and not as a tool for the training or teaching of humans.
In addition, although a majority of medical information is kept untouched due to the medical NER model in the Masker, some medical entities that are classified as non-critical may be masked and replaced. As a result, and because it is really difficult to automate and expensive to manually evaluate, we are not able to guarantee that all generated records are medically sound. Therefore, we recommend using the synthetic data in cases where medical soundness is not mandatory.

### v2

@@ Abstract
In this paper, we present a system that generates synthetic free-text medical records, such as discharge summaries, admission notes and doctor correspondences, using Masked Language Modeling (MLM). Our system is designed to preserve the critical information of the records while introducing significant diversity and minimizing re-identification risk. The system incorporates a de-identification component that uses Philter to mask Protected Health Information (PHI), followed by a Medical Entity Recognition (NER) model to retain key medical information. We explore various masking ratios and mask-filling techniques to balance the trade-off between diversity and fidelity in the synthetic outputs without affecting overall readability. Our results demonstrate that the system can produce high-quality synthetic data with significant diversity while achieving a HIPAA-compliant PHI recall rate of 0.96 and a low re-identification risk of 0.035. Furthermore, downstream evaluations using an NER task reveal that the synthetic data can be effectively used to train models with performance comparable to those trained on real data. The flexibility of the system allows it to be adapted for specific use cases, making it a valuable tool for privacy-preserving data generation in medical research and healthcare applications.
@@ I Introduction
Among numerous synthetic data generation techniques, Casual Language Modeling (CLM) is the one currently recognized as the state-of-the-art. Encoder-decoder transformers and GPTs have showcased their capabilities to capture the statistical properties of real discharge summaries with moderate privacy risks. However, to our knowledge, no work has tried to use Masked Language Modeling for this task. In our previous work [12], we introduced the first system to generate synthetic data using Masked Language Modeling (MLM) and compared it to Casual Language Modeling at multiple synthetic generations tasks to conclude that MLM was able to outperform the latter at some synthetic data generation tasks while preserving CLM’s main ability to infer complex knowledge and relations.
@@
This paper aims to introduce a system specifically designed to generate synthetic free-text medical reports, including discharge summaries, admission notes and correspondences between doctors using Masked Language Modeling. This system includes a de-identification phase using a state-of-the-art de-identification tool called Philter [13], which does not rely on deep learning and is specifically designed for the recognition of protected health information (PHI). While most other works ignore de-identification or assume that input data are already manually labeled for de-identification, we aim to include de-identification directly within our system. Additionally, our system includes a Medical Entity Recognition model to select and retain critical medical information from the original medical records. The final system is capable of taking raw, neither annotated nor de-identified letters, and generates a large amount of unique synthetic letters that retain the critical medical information of the original letter and for which users can control its resemblance/diversity trade-off. Finally, it guarantees the reduction in the number of patient health identifiers to nearly zero at the cost of slightly reduced data fidelity. The system, training, and evaluation codes are made publicly available at https://github.com/SamySam0/SynDeidMLM.
@@ II Related Work
Using a similar system, [15] produced synthetic unstructured free-text medical data with low re-identification risk. However, the authors experienced high similarities between generated and original data, leading to reduced diversity in the synthetic data and limiting its potential in large-scale generations for uses such as oversampling. Additionally, they did not focus on de-identification and assumed that synthetic data generation reduces re-identification risk by creating new patient records with different content. As a result, they stated that synthetic data produced must undergo rigorous de-identification of PHI elements before they can be distributed for public use.
@@
In one of the latest works on medical synthetic data generation, [18] analyzed whether GPT-3.5 could generate high-quality discharge summaries. They observed that GPT-3.5 was likely to reproduce and list all concepts mentioned in the prompt. In addition, GPT-3.5 generated unnatural texts which missed necessary supporting information and tended to introduce spurious information. These are due to the hard controllability of autoregressive CLM models as to how much is kept or how much is removed, resulting in a lack of diversity in the synthetic data. That was indeed validated by clinician-evaluators who noted “the correctness of the generated discharge summaries while suffering in variety, supporting information, and coherence in narratives” (hallucinations). Moreover, they revealed that GPT-3.5 failed to present diagnoses as interconnected events, which is something that may be improved through the selection of tokens to be masked when using MLM approaches.
Finally, in one of our parallel works [19], we presented an alternative approach for generating synthetic de-identified discharge summaries, using a different dataset and system architecture. This alternative system includes rule-based components, offering a complementary perspective to the work described here. We highly recommend readers explore this paper for additional insights into synthetic medical data generation.
@@ III-A The Masker
Medical Entity Recognition. The second phase in the Masker is to detect medical entities that we believe are critical information which build the core of the letter and should not be masked to preserve the letter’s intention. For this purpose, we leverage a pre-trained Medical Named Entity Recognition (NER) model which was fine-tuned on i2b2-2010 by Stanza [^1] https://stanfordnlp.github.io/stanza/available_biomed_models.html. This NER extracts 3 categories of medical entities, namely PROBLEM, TEST, and TREATMENT; and reported a micro-averaged F1 score of 88.13% on the same dataset. We use this tool to tag and protect every remaining entity in our original letter that belongs to any of these medical categories, such that they can remain in the structure of the masked letter and guide the synthetic generation. Note that the NER can be replaced at the need to select different types of entities and critical information. Alternatively, a ratio can be set up to control how much of each medical entity’s category should be kept.
Part-of-Speech Tagger. The third and final phase in the Masker is to detect part of speech within the original letter. To this end, we use Stanza’s POS which assigns a part-of-speech to each remaining word based on both its definition and its context. Given the POS of all remaining entities, we select and mask a random subset of tagged entities in order to create diversity in the synthetic generations. To do so, we define a set of ratios, with values between 0 and 1 for each POS category, where values closer to 0 imply less masking and ones closer to 1 have higher masking rates. For example, we could give the set {NOUN: 0.7, VRB: 0.5} to approximately mask 70% of nouns, 50% of verbs and none of the other POS categories. We leave the choice of masking ratios to the user based on their own use cases. Results of experiments on different presets are given in section V.
@@ III-B The Mask-Filling System
MLM Model. The MLM model is a deep language model that takes in the masked letter and provides a probability distribution over all possible words to replace the masked entity. As our MLM model, we selected a fine-tuned instance of the famous pre-trained encoder BERT, named Bio_ClinicalBERT [^2] https://huggingface.co/emilyalsentzer/Bio_ClinicalBERT. This model was initialized from BioBERT[21] and fine-tuned on all notes from MIMIC III [22], a database containing electronic health records from ICU patients at the Beth Israel Hospital in Boston, MA. This makes it an excellent baseline trained on both biomedical and clinical data. We then further trained this model on our specific task using the 790 training letters provided by our dataset to obtain the final version of our MLM model. Please note that we did not try alternative models as a baseline for our system. However, we truly encourage further studies to experiment with different baseline models.
@@
  • Simultaneous Chunk Filling: The first filling method takes a chunk of the masked letter and passes it to the MLM model. The latter sends back the probabilities for each masked entity. For each masked entity, the Mask-Filling component selects which word to replace it with based on the probability distribution. The user can decide to pick the word with the highest probability deterministically, or, in order to promote diversity, he can choose to select each word stochastically from the probability distribution at the cost of a slightly reduced fidelity. Here, fidelity is affected as we do not select the most probable word anymore, but instead draw from the MLM’s normal distribution.
@@ IV-B Hyperparameter tuning and Training
During training, we perform a grid search to select the most optimal set of hyperparameters from the following values: $\alpha\in\{1\times 10^{-4},5\times 10^{-5},3\times 10^{-5}\}$, $\beta\in\{8,16\}$, $\phi\in\{0.75,1.0\}$ and $\psi\in\{0.30,0.50\}$; where $\alpha$ is the learning rate of the MLM model, $\beta$ is the training batch size, $\phi$ is the PHI’s masking proportion and $\psi$ is the overall masking probability. For convenience, we select the optimal number of training epochs through early stoppage with patience of $p=2$. While we agree that more advanced hyperparameter search methods exist, such as Bayesian Optimization [26] or Optuna [27], we decided to opt for grid search due to computational limitations.
We split the dataset into 80% training and 20% validation, using a random split. We once again recognize that k-fold cross-validation is usually more accurate, but are constrained by the same computational resources. For each possible set of hyperparameters, a new instance of the system is created. Then, during its training, training samples are re-processed at each epoch with random masking of up to $\psi$ percent, including $\phi$ percent of all PHI entities. This is made to allow the model to see varied versions of the same sample, increasing the diversity of cases it can learn from and thus reducing overfitting. In contrast, the validation set is masked consistently across all epochs to ensure fair comparison.
@@ IV-C System instances
  • System_S_0.7: Similarly to System_S_0.5, this instance masks all PHI entities and none of the medical entities captured by the NER. However, it masks 70% of NOUNS, VERBS and ADJECTIVES for increased diversity, and uses the same Simultaneous Chunk-filling algorithm.
@@ IV-D Evaluation Metrics
Data privacy evaluation is crucial when considering the sharing of synthetic medical data. As our current dataset has been labeled by multiple professionals following the official HIPAA-PHI de-identification rules, we evaluate the privacy level of our model by calculating the F1-score to how much of the PHIs were identified and replaced by our system according to the annotated data, and how much re-identification occurred on average post generation.
@@ V Results and Discussion
In this section, we present the results of all system instances at the three evaluation phases described in section IV. To read full synthetic samples, please refer to Appendix Examples of generated synthetic letters. Note that, as there is no existing work on the synthetic generation of de-identified medical records with the same dataset, we are limited to comparing different configurations of our system internally.
@@ V-A Lexical Similarity Evaluation against References
When looking specifically at the results of the System_S instances, increasing the masking ratios penalizes the similarity more than with the iterative systems. In fact, on the same ratios (0.7), all ROUGE scores are about 3 points lower and the BERTScore 0.32 lower. That showcases one of the advantages of Iterative Mask Filling being that each mask token to fill is surrounded by original or predicted tokens, which adds to the MLM context and reduces overall uncertainty. On the other hand, when increasing System_Is’ masking ratios from 0.7 to 0.9, the ROUGE scores decrease between 3 and 5 points, while the BERTScore only decreases by less than half a point. This shows that while the generated letter is lexically further away from the original letter, the meaning is mostly preserved.
Finally, Table II presents the overlap between the top 5, 10, 20, 50, and 100 features identified across the real and synthetic datasets, without stopwords, to further evaluate the distance of the synthetic data to the original one. The results confirm that increasing the masking ratios and utilizing the iterative filling algorithms further promotes diverse generations while conserving an acceptable amount of core information.
@@ V-C1 Downstream NER Task
Before starting, we split the testing set into training and testing subsets. Then the downstream pipeline is presented as follows. First, the original letters are passed through the system to generate synthetic letters. Then, the original letters and synthetic letters are passed through SciSpacy [^5] https://allenai.github.io/scispacy/ en_ner_bc5cdr_md, a SpaCy [^6] https://spacy.io/ NER model trained on the BC5CDR corpus to recognize DISEASE and CHEMICAL entities, which achieved an F1-score of 0.843 on their dataset. Note that there exist alternative NER models that could be used here, something that could be worth exploring in future works. After extraction, we obtain the detected medical entities from both the original and synthetic letters. These entities are then stored to act as an original and synthetic entities dataset to train a Spacy NER model from scratch. That is, we train a first version of the Spacy NER model on the medical NER data extracted from the original training subset, and a second version on the ones extracted from the synthetic training subset. Finally, we evaluate both versions of Spacy on the testing subset and compare their performance.
Moreover, to see how much improvement can be obtained by further augmenting the synthetic dataset, we repeat this experiment and generate two synthetic letters per original letter.
@@ V-C2 Results
The results of the downstream task are given in Table IV. As they demonstrate, all systems could achieve scores comparable to those obtained on the original letters. An interesting observation is that the more we increase the masking ratios, the higher the scores tend to be. That is probably due to the increased diversity generated in synthetic samples, allowing to generate more diverse samples for Spacy to train on. Additionally, it is interesting that the iterative mask-filling technique could achieve the highest score by a slight margin, which is in accordance with the lexical results obtained above.
@@
TABLE IV: Average Precision, Recall and F1-score for two labels (DISEASE and CHEMICAL) using Synthetic data ×$\times 1$, ×$\times 2$, and Real data on the testing dataset.
@@ V-D Data Privacy Evaluation
Secondly, we calculated the re-identification risk, i.e. the probability that our MLM model re-introduces a valid protected health information masked by the Masker. To evaluate the re-identification risk, we leverage the risk of presence disclosure which assesses an attacker’s ability to determine if any real patient records in their possession were used to train the MLM model by comparing these records against the synthetic patient dataset. Assessing the risk of presence disclosure ensures the privacy of individuals whose data was used to train a decision model, as well as the interests of the healthcare entity where the individual received treatment. Thus, presence disclosure is a widely evaluated measure of re-identification risk. As a result, we observed that our MLM model re-identified masked PHI entities of more than two tokens with a probability of only 0.035. To give further insights into the re-identification risk, we calculated the rate of longest common substrings with lengths of 3 or more, 5 or more, and 7 or more, between original and synthetic PHIs. The outcomes showed rates of only 0.098, 0.020 and 0.009, respectively.
@@ VI-A Limitations
Upon careful analysis of the generated samples, we observed that consistently filling temporal information throughout the synthetic letter, and aligning it with the original data, presents a significant challenge. Furthermore, other interconnected events such as two names within a discussion may be complicated to assign consistently when one is not present in the context’s window at the moment of generation. This is an issue that our system may encounter and is rather important as medical data inherently consists of time series, where interconnected and temporal information are critical for numerous applications, such as modeling the progression of diseases. As a result, we recommend using our system as a means to generate synthetic data used to enhance machine learning performance, and not as a tool for the training or teaching of humans.
In addition, although a majority of medical information is kept untouched due to the medical NER model in the Masker, some medical entities that are classified as non-critical may be masked and replaced. As a result, and because it is really difficult to automate and expensive to manually evaluate, we are not able to guarantee that all generated records are medically sound. Therefore, we recommend using the synthetic data in cases where medical soundness is not necessary.